OpenStudy (anonymous):

Find the slope of a line passing through the points (1980,19000) and (2000,4000) help PLEASE

OpenStudy (anonymous):

shoot typo, it should read (1980,19000) and (2000,41000) NOT (4000)

jimthompson5910 (jim_thompson5910):

Subtract the y coordinates: 41000 - 19000 = 22000

jimthompson5910 (jim_thompson5910):

Subtract the x coordinates (in the same order). Tell me what you get.

OpenStudy (anonymous):

20

OpenStudy (anonymous):

is the x coordinates?

jimthompson5910 (jim_thompson5910):

no, it's the change in y over change in x

jimthompson5910 (jim_thompson5910):

slope = (difference in y coordinates)/(difference in x coordinates) slope = (22000)/(20) slope = 1100

OpenStudy (anonymous):

OH! I see, and then I multiply 1100 by 20 which gave me 22,000 and add that to my 41000 which gives me a median family increase to 63000 in year 2005.

jimthompson5910 (jim_thompson5910):

No, you found the slope to be m = 1100

OpenStudy (anonymous):

i'm confused

jimthompson5910 (jim_thompson5910):

now use this in y = mx+b along with x = 1980 and y = 19000 to find b

OpenStudy (anonymous):

first year algebra, forgive me, can you help me walk through it without giving me the answer?

jimthompson5910 (jim_thompson5910):

y = mx+b y = 1100x + b 19000 = 1100*1980 + b Now solve for b

jimthompson5910 (jim_thompson5910):

Does that make sense?

OpenStudy (anonymous):

2178000 when i multiply 1100*1980.

jimthompson5910 (jim_thompson5910):

good so far, what's next?

OpenStudy (anonymous):

2159000 is b?

OpenStudy (anonymous):

think i'm doing it right

jimthompson5910 (jim_thompson5910):

it's negative though

jimthompson5910 (jim_thompson5910):

since 19000 - 2178000 = -2159000

jimthompson5910 (jim_thompson5910):

so y = mx+b becomes y = 1100x - 2159000

jimthompson5910 (jim_thompson5910):

so in 2005, the median income is... y = 1100x - 2159000 y = 1100(2005) - 2159000 Keep going to get the answer

OpenStudy (anonymous):

46500

OpenStudy (anonymous):

thanks Jim. Let me know how I did. :)

jimthompson5910 (jim_thompson5910):

that's it, you nailed it

OpenStudy (anonymous):

You rule! Thank you. :)

jimthompson5910 (jim_thompson5910):

you're welcome